OpenStudy (anonymous):

A line has a slope of –3 and a y-intercept of (0, –1). What is the equation of the line that is parallel to the given line and passes through the point (–3, 1)? A. y = –3x – 8 B. y = 3x + 10 C. y = –3x + 4 D.

OpenStudy (imstuck):

Do you know how to tell if 2 lines are parallel to each other?

OpenStudy (anonymous):

Not really

OpenStudy (imstuck):

They have the same slope. The slope of the first line is given to you as -3. That means that the second line you find an equation to will also have a slope of -3.

OpenStudy (imstuck):

Do you know how to use the point-slope formula?

OpenStudy (anonymous):

Yah it's y-y1=m(x-x1)

OpenStudy (imstuck):

Good! That's half the battle! It's easy from here on out!

OpenStudy (imstuck):

Use the slope they gave you (-3) in place of the "m" in your point-slope equation and then use the point they gave you (-3, 1) and replace the y1 with the 1 and the x1 with the -3. Like so:

OpenStudy (imstuck):

\[y-1=-3(x-(-3))\]or\[y-1=-3(x+3)\]

OpenStudy (imstuck):

Now "do the math". Distribute the -3 into the parenthesis and you get\[y-1=-3x-9\]

OpenStudy (imstuck):

Add the 1 to both sides to isolate the y:\[y=-3x-9+1\]which of course is the same thing as\[y=-3x-8\]And that is your new equation with the slope of -3 that goes through the point (-3, 1)

OpenStudy (imstuck):

It's A in your choices.

OpenStudy (anonymous):

Ok thank you so much

OpenStudy (imstuck):

You're very welcome! Just do all of those types the same way...they all "work" the same!